1/2x2/3x3/4x4/5x9/10


1
Expert's answer
2022-03-29T12:08:18-0400

"\frac{1}{2}*\frac{2}{3}*\frac{3}{4}*\frac{4}{5}*\frac{9}{10}=\frac{1}{5}*\frac{9}{10}=\frac{9}{50}=0.18."